# Kernel-U-Net: Symmetric and Hierarchical Architecture for Multivariate Time Series Forecasting

### v2

@@ A.3 Dataset and Experiment
ETTh1
96 0.053 0.177 0.055 0.179 0.053 0.177 0.056 0.18 0.079 0.215 0.071 0.206 0.283 0.468 0.211 0.379
192 0.069 0.205 0.071 0.215 0.069 0.204 0.071 0.204 0.104 0.245 0.114 0.262 0.234 0.409 0.228 0.403
336 0.076 0.221 0.076 0.22 0.081 0.226 0.098 0.244 0.119 0.27 0.107 0.258 0.386 0.546 0.179 0.355
720 0.077 0.224 0.087 0.232 0.08 0.226 0.189 0.359 0.127 0.28 0.126 0.283 0.475 0.629 0.260 0.444
ETTh2
96 0.114 0.267 0.129 0.282 0.129 0.278 0.131 0.279 0.128 0.271 0.153 0.306 0.217 0.379 0.240 0.398
192 0.146 0.306 0.168 0.328 0.169 0.324 0.176 0.329 0.185 0.33 0.204 0.351 0.281 0.429 0.270 0.429
336 0.157 0.321 0.171 0.336 0.194 0.355 0.209 0.367 0.231 0.378 0.246 0.389 0.293 0.437 0.263 0.422
720 0.182 0.347 0.223 0.38 0.225 0.381 0.276 0.426 0.278 0.42 0.268 0.409 0.218 0.387 0.264 0.423
ETTm1
96 0.026 0.122 0.026 0.121 0.026 0.122 0.028 0.123 0.033 0.14 0.056 0.183 0.049 0.171 0.211 0.393
192 0.039 0.15 0.039 0.15 0.039 0.149 0.045 0.156 0.058 0.186 0.081 0.216 0.157 0.317 0.262 0.42
336 0.051 0.173 0.053 0.173 0.052 0.172 0.061 0.182 0.071 0.209 0.076 0.218 0.289 0.459 0.424 0.559
720 0.068 0.202 0.073 0.206 0.073 0.207 0.08 0.21 0.102 0.25 0.11 0.267 0.43 0.579 0.413 0.552
ETTm2
96 0.062 0.182 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208 0.128 0.275
192 0.089 0.224 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275 0.155 0.308
336 0.114 0.259 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302 0.208 0.358
720 0.156 0.312 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321 0.261 0.405
Table 4: Comparison of 16 variants of Linear K-U-Net in baseline setting and without using Skip connection(-Skip), Random Erase(-RE), Chanel Independence(-CI). The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined for minimum score(Min) and average score(Mean).
Method Min (Base) Min (-Skip) Min (-RE) Min (-CI) Mean (Base) Mean (-Skip) Mean (-RE) Mean (-CI)
Metric MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE MSE Top5-
MMSE
ETTh1
96 0.368 0.368 0.375 0.375 0.370 0.369 0.459 0.451 0.376 0.374 0.393 0.385 0.377 0.373 0.488 0.483
192 0.404 0.402 0.412 0.413 0.402 0.402 0.506 0.495 0.409 0.407 0.423 0.423 0.411 0.407 0.551 0.538
336 0.423 0.421 0.443 0.441 0.427 0.422 0.535 0.519 0.443 0.430 0.489 0.455 0.442 0.431 0.581 0.549
720 0.457 0.439 0.461 0.453 0.461 0.438 0.629 0.576 0.492 0.472 0.531 0.515 0.488 0.470 0.690 0.598
ETTh2
96 0.274 0.272 0.273 0.273 0.274 0.273 0.423 0.397 0.302 0.287 0.299 0.290 0.300 0.286 0.458 0.415
192 0.336 0.353 0.337 0.355 0.336 0.354 0.414 0.442 0.370 0.375 0.376 0.382 0.372 0.374 0.445 0.474
336 0.361 0.361 0.364 0.364 0.362 0.361 0.412 0.422 0.402 0.384 0.404 0.392 0.399 0.384 0.428 0.444
720 0.398 0.394 0.397 0.397 0.395 0.394 0.453 0.456 0.451 0.418 0.459 0.436 0.444 0.420 0.468 0.519
ETTm1
96 0.288 0.284 0.289 0.290 0.288 0.284 0.321 0.318 0.296 0.292 0.303 0.298 0.296 0.291 0.346 0.333
192 0.324 0.327 0.325 0.324 0.326 0.324 0.359 0.358 0.335 0.331 0.341 0.336 0.334 0.331 0.383 0.373
336 0.359 0.358 0.353 0.354 0.359 0.357 0.391 0.391 0.366 0.363 0.374 0.368 0.366 0.362 0.418 0.406
720 0.408 0.404 0.403 0.403 0.409 0.403 0.452 0.447 0.418 0.411 0.419 0.413 0.418 0.410 0.478 0.467
ETTm2
96 0.162 0.162 0.162 0.161 0.163 0.161 0.213 0.204 0.168 0.165 0.170 0.165 0.168 0.164 0.228 0.210
192 0.218 0.217 0.216 0.215 0.217 0.216 0.298 0.274 0.228 0.223 0.229 0.224 0.228 0.222 0.318 0.286
336 0.270 0.268 0.268 0.267 0.265 0.265 0.372 0.345 0.283 0.277 0.281 0.277 0.284 0.272 0.394 0.356
720 0.342 0.349 0.343 0.347 0.343 0.347 0.410 0.408 0.354 0.361 0.351 0.361 0.364 0.356 0.434 0.421
Average decrease +0.42% +0.77% +0.03% -0.20% +24.19% +21.45% +2.24% +2.30% +0.03% -0.32% +25.92% +23.72%
Experiment details
@@
We use a 4-layer K-U-Net for experiments. The list of multiples are respectively [4, 4, 3, 7] and [4, 6, 6, 5] for look-back windows $L$ and forecasting horizons $T$ in {336, 720}. Remark that $T$ is fixed to be 336 for shorter forecasting cases. The bottom patch length is 4 and its width is 1 as we follow the Channel-Independent (CI) setting in Zeng et al. [2023]. In case of not using CI, the patch width equals to feature size. The hidden dimensions are 128 for multivariate and univariate tasks in general, except that they are 256 on the Traffic dataset. The learning rate is selected in $[0.00005,0.001]$ depending on the dataset. The training epoch is 50 and the patience of early stopping is 10 in general.
We apply weighted early stopping where the triggering MSE (TMSE) is computed with $TMSE_{e+1}=\tau*TMSE_{e}+(1-\tau)*VMSE_{e}$ , where $\tau$ is a hyper-parameter to be set, $e$ is the index of the episode, and VMSE is the running MSE on a validation set. The parameter $\tau$ sets the sensitivity of early stopping and is empirically selected. In general, it is set to be 0.9 in the variants searching stage and 0.5 in the fine-tuning stage.
The MSE loss function is defined with $:\mathcal{L}=\mathbb{E}_{\boldsymbol{x}}\frac{1}{B*M}\sum_{i=1}^{B*M}\left\|\hat{\boldsymbol{x}}_{L+1:L+T}^{(i)}-\boldsymbol{x}_{L+1:L+T}^{(i)}\right\|_{2}^{2}$ , where $B$ is the batch size and $M$ is the multiple of feature size unit of time series.
A.4 Search of model variants
We searched 30 variants of K-U-Net across Linear_[$0$,…,$1$] and {LSTM, Transformer}_[0100,…,$1$] for the ETT, Weather datasets and 3 subsets. The lookback window is typically 720, and we additionally tested a version of 336 on the ETT dataset. The result is reported with the average of running Top5 Minimum MSE (Top5MMSE) on the validation and test set in a single run. This measures the extreme performance of a variant and its potential in the fine-tuning stage. We use relative score (RS) to locate best-performed variants (i.g. $RS=1$ means that it’s the best among all candidates). We plot the RS of variants of 5 datasets in Figures 9 and Figures 10.
Based on the search result, We note that Linear and Multilayer linear kernels are more adapted in the ETT dataset than our Transformer and LSTM kernels. We remark that K-U-Net Linear_1000 is outstanding in the Electricity (S) dataset and thus selected for full examination.
A.5 Result of selected variants
In Table 5 6 7, we note selected models that are examined prudently with 5 runs in a fine-tuned context and measures with average MSE and MAE. We include the Table 9 showing that K-U-Net has a small standard deviation error across 5 random runs.
A.6 Result of univariate time series forecasting

### v3

@@ III-B2 Hierarchical processing with kernels
ETTh1 96 0.355 0.388 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713 0.985 0.74
192 0.388 0.412 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792 1.17 0.855
336 0.407 0.427 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809 1.208 0.886
720 0.430 0.454 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865 1.34 0.899
ETTh2 96 0.269 0.335 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525 1.335 0.936
336 0.355 0.400 0.329 0.384 0.357 0.400 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835 1.444 0.96
720 0.384 0.435 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625 3.498 1.631
ETTm1 96 0.275 0.331 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571 0.849 0.669
192 0.320 0.361 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669 0.928 0.724
336 0.349 0.380 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871 1.058 0.786
720 0.401 0.412 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823 0.955 0.703
ETTm2 96 0.157 0.243 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453 0.487 0.529
192 0.213 0.283 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563 0.789 0.705
336 0.266 0.320 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 1.256 0.904
720 0.343 0.377 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338 2.698 1.297
Electricity 96 0.128 0.219 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368 - -
192 0.145 0.234 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 - -
336 0.160 0.250 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394 - -
720 0.196 0.283 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 - -
Traffic 96 0.354 0.229 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391 - -
192 0.372 0.262 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 - -
336 0.388 0.270 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42 - -
720 0.430 0.269 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472 - -
Weather 96 0.142 0.183 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384 - -
192 0.187 0.226 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544 - -
336 0.238 0.269 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523 - -
720 0.308 0.323 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 - -
At the stage of the decoder, the operations are reversed. We start with $\mathcal{Q}^{(n+m)}_{in}=\mathcal{Z}$ of shape $(B,1,D_{h})$, a set of input patches to the decoder. We send it to decoder kernel and get an output $\mathcal{Q}^{(n+m)}_{out}$ in shape $(B,M_{m},D_{h})$, then we reshape it to $\mathcal{Q}^{(n+m-1)}_{in}$ in shape $(B\cdot M_{m},1,D_{h})$ for the next kernel. At the end of this iteration over multiples $,\{M_{m-1},$...,… , $M_{2},$ $L_{n},$ ,...,… , $L_{2},$ $L_{1}\}$, we finally have a set of patches $\mathcal{Q}^{(1)}_{out}$ in the shape $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k},L_{1},M_{1})$. The last operations are reshaping it into $(B,\prod_{2}^{m}M_{k},\prod_{2}^{n}L_{k},L_{1},1,M_{1})$ , transposing it into $(B,1,\prod_{2}^{n}L_{k},L_{1},\prod_{2}^{m}M_{k},M_{1})$ and reshaping it into $(B,L,M)$ as final output.
III-B3 Kernel Wrapper
The Kernel Wrapper (KW) requires necessary parameters such as kernel $\phi$, input patches $X$, output patches $Z$, patches set size $\mathcal{B}$, input patch length $J_{in}$ and dimension $D_{in}$, output patch length $J_{out}$ and dimension $D_{out}$. The kernel wrapper initiates an instance of a given kernel and calls it in the forward function. It reshapes the input patches, executes with the kernel inside, and then checks the output shape. The wrapper processes $X$ and outputs $Z$ in the encoder, or it processes the sum of $X$ and an encoder output via skip connection, then gives an output in the decoder (Algorithm B).
III-B4 Formulation of Kernel Operation
We add enc and dec in the index of variables to differentiate their utilization in the encoder and decoder. We denote $\mathcal{P}=X_{enc}$ and $\mathcal{Q}=X_{dec}$ the set of patches, $l\in\{1,...,n,n+2,...,n+m\}$ the index of layers, and $\mathcal{B}$ the patches set size. By default, an encoder kernel at layer $l$ receives $\mathcal{P}^{(l)}_{in}\in\mathbb{R}^{\mathcal{B}^{(l)}\times J^{(l)}_{enc\_in}\times D^{(l)}_{enc\_in}}$ and gives $\mathcal{P}^{(l)}_{out}\in\mathbb{R}^{\mathcal{B}^{(l)}\times 1\times D^{(l)}_{enc\_out}}$. The decoder kernel at layer $l$ receives $\mathcal{Q}^{(l)}_{in}\in\mathbb{R}^{\mathcal{B}^{(l)}\times 1\times D^{(l)}_{dec\_in}}$ and outputs $\mathcal{Q}^{(l)}_{out}\in\mathbb{R}^{\mathcal{B}^{(l)}\times J^{(l)}_{dec\_out}\times D^{(l)}_{dec\_out}}$.
We recall that $i^{(l)},j^{(l)}$ are the indices for multiples of length and feature at layer $l$. Given a set of input $\mathcal{P}^{(l)}_{in}$ at layer $l$, the output $\mathcal{P}^{(l)}_{out}$ of kernel $\phi^{(l)}_{enc}$ in encoder is written :
$\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}}=\phi^{(l)}_{enc}(\mathcal{P}^{(l)}_{in,i^{(l)},j^{(l)}})$ (2)
The decoder kernel $\phi^{(l)}_{dec}$ at layer $l$ takes the sum of input $\mathcal{Q}^{(l)}_{in}$ and encoder output $\mathcal{P}^{(l)}_{out}$ via the skip connection as input and produces $\mathcal{Q}^{(l)}_{out}$ as output:
$\mathcal{Q}^{(l)}_{out,i^{(l)},j^{(l)}}=\phi^{(l)}_{dec}(\mathcal{Q}^{(l)}_{in,i^{(l)},j^{(l)}}+\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}})$ (3)
Remark that in case of $l=n+m$, we have $\mathcal{Q}^{(l)}_{in,i^{(l)},j^{(l)}}+\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}}=\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}}$ because there are no higher layers and the kernel only process the encoder output.
III-B5 Creation of Kernel U-Net
We create the encoder, decoder, and the U-Net in order. The algorithm passes parameters describing the multiples, kernels, and hidden dimensions. Let us note the input length $L$, feature dimension $M$, concatenated lists of multiples of look-back window and feature $,\{L_{2},$...,… , $L_{n},$ $M_{2},$ ,...,… , $M_{m}\}$, list of hidden dimension of same size $,\{D_{h},$...,… , $D_{h}\}$, patch size $L_{1}$ and feature unit $M_{1}$ , a list of kernels $\{\phi^{(1)}_{enc},$ $\phi^{(2)}_{enc},$ ,...,… , $\phi^{(n)}_{enc},$ $\phi^{(n+2)}_{enc},$ ,...,… , $\phi^{(n+m)}_{enc}\}$, latent vector length $J_{h}$ and dimension $D_{h}$. We also use next and prev to iterate over the index $l$. We set the hidden dimension of intermediate output vectors within layers to be equal to that of latent vector for simpler description. It corresponds to channel size in a convolutional network and can be augmented for a larger passage of information if necessary. We describe the creation of the K-U-Net encoder in Algorithm A.
The decoder is symmetrical to the encoder and applies kernels in reverse order. More precisely, the decoder takes $J_{h}$ and $D_{h}$, multiples $J\in\{L_{1},$ $L_{2},...,$ $L_{n},$ $M_{2},$ ,...,… , $M_{m}\}$, kernels $,\{\phi^{(n+m)}_{dec},$...,… , $\phi^{(n+2)}_{dec},$ $\phi^{(n)}_{dec},$ ,...,… , $\phi^{(2)}_{dec},$ $\phi^{(1)}_{dec}\}$ and initiates kernel wrappers KW($\phi^{(n+m)}_{dec},$ $J_{h},$ $D_{h},$ $J_{n+m},D_{n+m})$, KW($\phi^{(l)}_{dec},$ $1,$ $D_{l\_prev},$ $J_{l},D_{l})$, KW($\phi^{(1)}_{dec},$ $1,$ $D_{2},$ $L_{1},M_{1})$ at the highest, intermediate, and lowest layers respectively.
The K-U-Net initiates an encoder and a decoder. In the forward function, the encoder processes the input and generates a list of outputs at each layer and a latent vector. It assigns the skip_out from encoder kernels to skip_in in decoder kernels. Then the decoder takes the outputs from the encoder via skip-connection and the latent vector to generate the final result.
Figure 2: Structure of Linear Kernel. a) Linear Kernel, b) Multi-layer perceptron (MLP) Kernel with Tanh activation, c) LSTM Kernel, d) Transformer Kernel
III-C Custom Kernels
III-C1 Linear kernel
The linear kernel is a simple matrix multiplication. Given $X\in\mathbb{R}^{\mathcal{B}\times J_{in}\times D_{in}}$, linear kernel $\phi$ reshape it to $(\mathcal{B},J_{in}\cdot D_{in})$ and process it as follow:
$Z=\phi(X)=XW+b$ (4)
where $Z$ is output of kernel, $W\in\mathbb{R}^{J_{in}\cdot D_{in}\times J_{out}\cdot D_{out}}$ is weight matrix and $b\in\mathbb{R}^{J_{out}\cdot D_{out}}$ is bias vector. Remark that the number of parameters of $W$ is $J_{in}\cdot D_{in}\cdot J_{out}\cdot D_{out}$ and this kernel operation is equivalent to the process with a 1D convolutional layer whose kernel size is $J_{in}$ or $J_{out}$.
III-C2 Multi-Layer Perceptron kernel
The multi-layer perceptron (MLP) kernel has an additional hidden layer and a non-linear activation function Tanh. The formulation is:
$Z=\phi(X)=Tanh(XW_{1}+b_{1})W_{2}+b_{2}$ (5)
where $Z$ is output of kernel, $W_{1}\in\mathbb{R}^{J_{in}\cdot D_{in}\times J^{\prime}\cdot D^{\prime}}$ and $W_{2}\in\mathbb{R}^{J^{\prime}\cdot D^{\prime}\times J_{out}\cdot D_{out}}$ are weight matrices , $b_{1}\in\mathbb{R}^{J^{\prime}\cdot D^{\prime}}$ and $b_{2}\in\mathbb{R}^{J_{out}\cdot D_{out}}$ are bias vectors, and $J^{\prime}=\frac{1}{2}(J_{in}+J_{out}),D^{\prime}=\frac{1}{2}(D_{in}+D_{out})$.
III-C3 Transformer kernel
The Vanilla Transformer[16] is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with ReLU activation. The transformer kernel in this work follows the classical structure (Figure 2).
III-C4 LSTM kernel
The LSTM kernel contains a classic LSTM cell[31] and a linear layer for skip connection. The hidden states of all time steps are combined with a linear layer in the next (Figure 2).
III-D Complexity Analysis
Since the K-U-Net is symmetric, we only study the complexity of the encoder layer and assume that the feature size is $1$. Let us suppose that a kernel is receiving a sequence of length $L$ where $L=\prod_{1}^{n}L_{i}$ and all $L_{i}$ are equal. As the patch size in the first layer of Kernel U-Net encoder is $L_{1}$, a kernel module will process $\frac{L}{L_{1}}$ patches of size $L_{1}$. Therefore, the complexity is $O(\frac{L}{L_{1}^{n}}\cdot g(L_{1}))$ where $g(L_{1})$ is the complexity inside the kernel in the function of patch size $L_{1}$ and $n$ is the index of the layer. In the case of using the linear kernel at the first layer, the complexity is $O(\frac{L}{L_{1}}\cdot L_{1})$= $O(L)$. In the case of using a classic transformer kernel, the complexity is $O(\frac{L}{L_{1}}\cdot L_{1}^{2})$= $O(L\cdot L_{1})$. Let us set $L_{1}=log(L)$, the complexity of the application of such a quadratic calculation kernel is $O(Llog(L))$. Moreover, if we apply the transformer kernel starting from the second layer, the complexity is dramatically reduced to $O(\frac{L}{L_{1}\cdot L_{2}}\cdot L_{2}^{2})$= $O(L)$. Following the same demonstration, the complexity of using LSTM kernels and MLP starting from the second layer is also bounded by $O(L)$.
IV Experiments and Results
IV-1 Datasets
We conducted experiments with our proposed Kernel U-Net on 7 public datasets, including 4 ETT (Electricity Transformer Temperature) datasets [^2] https://github.com/zhouhaoyi/ETDataset (ETTh1, ETTh2, ETTm1, ETTm2), Weather [^3] https://www.bgc-jena.mpg.de/wetter/, Traffic [^4] http://pems.dot.ca.gov and Electricity [^5] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014. These datasets have been benchmarked and publicly available on[2] and the description of the dataset is available in [12].
Here, we followed the experiment setting in [12] and partitioned the data to $[12,4,4]$ months for training, validation, and testing respectively for the ETT dataset. The data is split to $[0.7,0.1,0.2]$ for training, validation, and testing for the Weather, Traffic, and Electricity datasets.
IV-2 Baselines and Experimental Settings
We follow the experiment setting in NLinear [12] and take $L$ step historical data as input then forecast $T$ step value in the future where $L\in\mathcal{L}=\{336,720\}$ and $T\in\mathcal{T}=\{96,192,336,720\}$. We replace the last value normalization by mean value normalization for ETT, Electricity, and Weather datasets, and apply instance normalization[32] for the Traffic dataset. We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) for evaluation as mentioned in [2].
We include recent methods: PatchTST[13], NLinear, DLinear[12], FEDformer [11], Autoformer [2], Informer [3], LogTrans [10] Yformer[15]. We merge the result reported in [13] and [12] for taking their best in a supervised setting and execute Yformer with default parameters in its Github [^6] https://github.com/18kiran12/Yformer-Time-Series-Forecasting.
IV-3 Experiment details
@@
We use a 4-layer K-U-Net for experiments. The list of multiples are respectively {4, 3, 7} and {6, 6, 5} for look-back windows $L\in\{336,720\}$. The bottom patch length is 4 and its width is 1. We reshape the input into $(B\cdot M,L,1)$ for processing with K-U-Net as we follow the channel-independent setting [12] and then reshape it back. The hidden dimensions are 128 for multivariate tasks. The learning rate is selected in $[0.00005,0.001]$. The training epoch is 50 and the patience of early stopping is 10 in general. We use MAE as the loss function.
IV-4 Model Variants Search
We propose 4 kernels for experiments with K-U-Net on 7 datasets. By replacing a linear kernel at different layers with other types of kernels, we search for variants that adapt the dataset. As there are too many variants to name, we note them as ”KUN_$<$kernel$>$_$<$replace_layer$>$ ($<$look-back_window$>$)”. For example, KUN_Linear_$\_0000$ $(720)$ means that the model is made of linear kernels at all layers and the look-back window size is 720, KUN_Transf_$\_0100$ means that a transformer kernel replaces the linear kernel at the second layer, KUN_Linear_$\_0110$ means that multilayer perceptron kernels replace the kernels at the second and third layers. We have composed 16 variants with MLP kernels and 7 variants with Transformer and LSTM kernels.
Figure 3: Search result of K-U-Net with Linear, MLP, LSTM, Transformer kernels on ETTh1 dataset.
Figure 4: GPU consumation of K-U-Net, PatchTST and Yformer.
$\text{RS}(V)=\text{Min}_{T\in\mathcal{T}}\frac{\text{Top5MMSE}(V,T)}{\text{Min}_{\{U\in\mathcal{V},L\in\mathcal{L}\}}\text{Top5MMSE}(U,T,L)}$ (6)
, where $U$, $V$ are variant models in the models set $\mathcal{V}$, $T$ is a forecasting horizon in $\mathcal{T}$, $L$ is a look back window in $\mathcal{L}$. Relative score notes the best-performed model with 1 and thus helps to identify the high-potential candidates for further fine-tuning examinations.
We observe in Figure 3 that the best variant for ETTh1 dataset is KUN_Linear_$\_0010$. Comparing the relative score of the K-U-Net of replace__\__layer code ($0010$, $0011$), ($0110$, $0111$) and ($0100$, $0101$), we remark that replacing the highest layer with Transformer and LSTM kernel degrades the performance because of overfitting. Furthermore, We observe in Figure 5 that the best variants for the Weather dataset are KUN_Linear_$\_0110$, KUN_LSTM_$\_0100$ and KUN_Transf_$\_0100$. Comparing the relative score of the K-U-Net of replace__\__layer code ($0010$, $0110$), ($0011$, $0111$) and ($0001$, $0101$) we remark that replacing the second layer with Transformer and LSTM kernel gains the performance for their expressiveness.
Among candidates in the search phase, we empirically choose 3 variants for fine-tuning experiments with 5 runs. The final result shows that the performance of Kernel U-Net exceeds or meets the state-of-the-art methods in multivariate settings in most cases.
IV-5 Multivariate time series forecasting result